OpenStudy (chrispaul03):

Fan And Medal -7+4n+9 how do I solve this?

alones (alones):

−7+4n+9 Combine Like Terms: =−7+4n+9 =(4n)+(−7+9) =4n+2 Answer: =4n+2

OpenStudy (anonymous):